Mitch practices playing the violin every evening for 1 hour and 45 minutes. What is the positive difference between the total degrees traveled by the minute hand and the total degrees traveled by the hour hand during Mitch’s practice? Express your answer as a decimal to the nearest tenth.
Answer by ankor@dixie-net.com(22740)

Mitch practices playing the violin every evening for 1 hour and 45 minutes.
What is the positive difference between the total degrees traveled by the minute
hand and the total degrees traveled by the hour hand during Mitch’s practice?
Express your answer as a decimal to the nearest tenth.
:
Change 1 hr 45 min to 1.75 hrs
Each hr the minute hand travels 360 degrees
1.75 * 360 = 630 degrees traveled by the minute hand
:
The hr hand travels 360/12 = 30 degrees each hr
1.75(30) = 52.5 degrees traveled by the hr hand
:
630 - 52.5 = 577.5 degrees